Am I doing this correctly? Can someone please post the answers? This Question: 15 pts This Quiz: 70 pts possib 24 of 27 (27 complete) Unger Autoparts Inc. issued $190,000 of 2%, 10-year bon
Answer #1
1. Effective-interest amortization table for the bonds through the first three interest payments:-
Effective Interest method of Bond Amortization:-
Unger Autoparts
Amortization table
Semi-annual interest Date Interest Payment    (1% of Maturity Value) Interest Expense to be recorded (2.5% of prreceding Bond carryig Amount) Bond Discount Amortization (Payment - Interest Expense) Bond discount Account Balance End of period Bond Carrying value (Beginning balance+ amortization)
A B C = B-A D = Precedind D -C E = $190000-D
Jan 31, 2017 $                       20,900 $           169,100
July 31, 2017 $            1,900 $                    4,228 $                                      2,328 $                       18,573 $           171,428
Jan 31, 2018 $            1,900 $                    4,286 $                                      2,386 $                       16,187 $           173,813
July 31, 2018 $            1,900 $                    4,345 $                                      2,445 $                       13,741 $           176,259
2. Journal entries:-
Date Account Titles and Explanation Debit Credit
Jan 31, 2017 Cash ($190000*89/100) $                                  169,100
Discount on bonds payable $                                    20,900
              Bonds Payable $                     190,000
(To record issue of bonds on discount)
July 31, 2017 Interest Expense $                                      4,228
              Interest Payable $                         1,900
              Discount on bonds payable $                         2,328
(To record the accrual of interest and amortization of discount)

Feel free to ask any clarification, if required. Please provide feedback by thumbs up, if satisfied. It will be highly appreciated. Thank you.